# RoadBEV: Road Surface Reconstruction in Bird’s Eye View

### v1

@@ Abstract
Road surface conditions, especially geometry profiles, enormously affect driving performance of autonomous vehicles. Vision-based online road reconstruction promisingly captures road information in advance. Existing solutions like monocular depth estimation and stereo matching suffer from modest performance. The recent technique of Bird’s-Eye-View (BEV) perception provides immense potential to more reliable and accurate reconstruction. This paper uniformly proposes two simple yet effective models for road elevation reconstruction in BEV named RoadBEV-mono and RoadBEV-stereo, which estimate road elevation with monocular and stereo images, respectively. The former directly fits elevation values based on voxel features queried from image view, while the latter efficiently recognizes road elevation patterns based on BEV volume representing discrepancy between left and right voxel features. Insightful analyses reveal their consistence and difference with perspective view. Experiments on real-world dataset verify the models’ effectiveness and superiority. Elevation errors of RoadBEV-mono and RoadBEV-stereo achieve 1.83cm and 0.56cm, respectively. The estimation performance improves by 50% in BEV based on monocular image. Our models are promising for practical applications, providing valuable references for vision-based BEV perception in autonomous driving. The code is released at https://github.com/ztsrxh/RoadBEV
@@ I Introduction
In recent years, the rapid development of unmanned ground vehicles (UGVs) has posed higher requirements for on-board perception systems. Real-time understanding of driving environment and conditions is vital for accurate motion planning and control [1, 2, 3]. For vehicles, road is the only contacting media with the physical world. Road surface conditions determine many vehicle characteristics and driving performance [4]. Road unevenness like bumps and potholes (as shown in Fig. 1 (a)) exacerbates ride experience of passenger vehicles, which is intuitively perceptible. Real-time road surface condition perception, especially the geometry elevation, immensely contributes to enhance ride comfort [5, 6].
@@ IV-C RoadBEV-stereo
Fig. 7 shows the architecture of RoadBEV-stereo. Similar with RoadBEV-mono and general stereo matching models, the left and right images first undergo feature extraction with shared weights. Settings about the CNN backbone are the same as RoadBEV-mono, i.e., EfficientNet-B6. The only difference is that the feature map resolution is interpolated to $\frac{1}{2}$ in the FPN. Explanations and analysis about resolution will be given below. The feature voxels are projected onto both the two image planes and query features from left and right perspectives. The left and right voxel features, $\boldsymbol{F}^{\prime}_{vox,l}$ and $\boldsymbol{F}^{\prime}_{vox,r}$, with the same shape $[C,N_{z},N_{y},N_{x}]$ are thereby obtained. In the architecture of stereo matching models, cost volume encoding similarities of the two feature maps is then built by correlation operation. We follow this paradigm but with subtraction to emphasize the discrepancy of left and right voxel features:
$\boldsymbol{V}_{discr}=\boldsymbol{F}^{\prime}_{vox,l}-\boldsymbol{F}^{\prime}_{vox,r}$ (2)
where $\boldsymbol{V}_{discr}$ is the built voxel discrepancy volume in BEV. The volume construction here also differs from stereo matching in terms of feature inputs. The vertical dimension of the feature voxel represents elevation candidates or proposals, which is equivalent to the disparity dimension in stereo matching. For stereo matching, input features are 3D and the disparity dimension is introduced during volume construction. Instead, our proposed method directly takes 4D voxel features as inputs whose proposal dimensions are introduced before volume construction.
Similar with RoadBEV-mono, the 4D BEV volume can be reshaped as a 3D tensor and processed with 2D convolutions, which efficiently saves computation. Nevertheless, to demonstrate its consistence with stereo matching models, we still treat it as a 4D volume and aggregate by 3D convolutions. Specifically, we stack six 3D-conv layers and three hourglass modules with $\frac{1}{2}$ resolution reduction. The channel dimension of BEV volume is finally reduced to 1. Since the number of vertical voxels $N_{z}$ may be different from the number of elevation classes $N_{c}$, the elevation proposal dimension is interpolated to have $N_{c}$ channels. The elevation feature map $\boldsymbol{F}^{\prime}_{ele}\in\mathbb{R}^{N_{c}\times N_{y}\times N_{x}}$ is thereby derived. The following output layers are the same with RoadBEV-mono.
Figure 7: Architecture of RoadBEV-stereo. The voxel defined under the left camera’s coordinate queries pixel features of the left and right feature maps. We construct discrepancy volume in BEV by subtraction between left and right voxel features. 3D convolutions then aggregate the 4D volume in BEV.
@@
Despite the benefits, requirements for accurate RSR with stereo images in BEV is more strict. First, high-resolution feature map is necessary to cooperate with the precise but narrow search range. Road image features are similar in a certain area as road texture are repetitive without regular patterns. High-resolution feature is therefore vital for identifying slight feature variation. This is also the reason why we leverage image feature maps with $\frac{1}{2}$ resolution. Then, effective similarity measurement that emphasizes feature discrepancy is crucial. The commonly adopted correlation operation (i.e., point-wise multiplication) in stereo matching neglects the magnitude difference and may introduce a holistic bias. Consequently, we propose constructing discrepancy volume by subtraction, which is more direct than correlation and magnify minor difference. Ablation studies in Section V-E further demonstrate their influence.
@@ V-A Implementation Details
We utilize the following metrics to evaluate model performance: absolute error (abs. err.) and root mean squared error (RMSE). We implement the models on PyTorch platform and train on RTX 3090 GPUs. Batch size is 8 for all experiments. Learning rate is set as 8e-4 for RoadBEV-mono, while 5e-4 for RoadBEV-stereo. The optimizer is AdamW with weight decay 1e-4. The OneCycle learning rate scheduler with linear decreasing strategy is used. The images are cropped to 960*528 to meet the size requirements. The RoadBEV-mono is trained for 50 epochs, while 40 for RoadBEV-stereo. Pre-trained weights on ImageNet is loaded. The compared models in Table I are also trained for the same epochs, with their default configurations in the provided codes.
@@ V-B Performance and Comparison
Method Abs. err. (cm) RMSE (cm)
Mono LapDepth[44] 2.81 3.12
PixelFormer[45] 2.65 2.86
iDisc[46] 2.64 2.88
AdaBins[43] 2.59 2.79
RoadBEV(Ours) 1.83 2.07
Stereo IGEV-Stereo[47] 0.651 0.797
PSMNet[48] 0.654 0.785
CFNet[49] 0.647 0.760
ACVNet[50] 0.596 0.723
GwcNet[51] 0.588 0.711
RoadBEV (Ours) 0.563 0.697
Further, all the stereo-based models in Table I outperform the monocular-based with immense margins. Our RoadBEV-stereo also performs better than the others with noteworthy improvements. The proposed discrepancy volume and estimation head jointly function and efficiently recognize feature patterns in BEV. The constraints on search range simplify similarity measurement and identification with effectiveness. The error level of 5.6mm covers almost all road unevenness causing vertical vibration to vehicles. Our model has great potential in on-board RSR for aiding planning and control systems of UGVs. Comparing the architectures of RoadBEV-mono (Fig. 5) and RoadBEV-stereo (Fig. 7), the principal promotion of RoadBEV-stereo lays in the utilization of extra information from another perspective (i.e., stereo image). The introduction of another view contributes to more than twice the performance enhancement. Therefore, stereo-based RSR is undoubtedly more promising and reliable than the monocular-based.
For more insights, we further visualize the distance-wise absolute error, as shown in Fig. 10. The whole grid is split into 15 segments along the longitudinal direction, with each segment representing a region of 33cm long. Absolute elevation errors of grids inside the segments are averaged. Instead of a global but coarse metric, this analysis method dives into model performance and gives more comprehensive evaluation. Our RoadBEV-mono thoroughly outperforms the compared depth estimation models among the whole range with a significant margin. The superiority is more prominent at far distance comparing with LapDepth. As analyzed in Section IV-B. The AdaBins performs well as it defines the searching bins rather than direct fitting without effective constraints.
@@ V-D Ablation Studies for RoadBEV-mono
We first explore the influence of class interval on model performance, as listed in Table II. We set different class intervals between 0.4$\sim$2cm, among which the metrics reach highest at 0.5cm. Results also show that metric difference for class resolutions $\{0.5,0.8,1.0,1.6\}$ is slight. Potential reason is that we utilize soft argmin to regress elevation instead of picking the peak class. The weighted sum operation compensate accuracy decrease caused by discrete class interval. Over-high resolution at 0.4cm results in $Nc$=100, which is too complicated to distinguish minute distinctness. Over-low resolution at 2.0cm is too sparse to accurately represent the continuous elevation. Class interval is thereby set as 0.5cm for both the proposed models.
@@
Class interval (cm) # classes Abs. err. (cm) RMSE (cm)
2.0 20 2.09 2.27
1.6 25 1.94 2.18
1.0 40 1.91 2.16
0.8 50 1.84 2.07
0.5 80 1.83 2.07
0.4 100 1.95 2.21
@@ V-E Ablation Studies for RoadBEV-stereo
As stated in Section IV-C, influence of volume construction is investigated. The first three experiments in Table IV exploit correlation to construct BEV volume. The metrics unexpectedly drop to quite low level regardless of other parameters, which are the approximate performance achieved by RoadBEV-mono. It is evident that point-wise multiplication cannot provide effective similarity measurement for the left and right feature queries. The estimation head suffers from identifying actual elevation among almost identical similarities. RoadBEV-stereo with correlation BEV volume actually reduces to RoadBEV-mono. Feature discrepancy by subtraction, rather than correlation by multiplication, is verified to be vital for exerting benefits brought by stereo information.
We further compare different settings of image and vertical voxel resolutions. Results imply that for a certain image resolution, optimal performance is achieved at 1.0cm voxel resolution. This is consistent with the conclusion drawn from RoadBEV-mono. It is also shown that for the same voxel resolution, the model with $\frac{1}{2}$ feature resolution is superior. For instance, the absolute error reduces from 0.730cm to 0.584cm (i.e., 20% improvement) when voxel resolution is 0.5cm. Another valuable discovery is that performance also collapses for 2.0cm voxel resolution. Sparse feature queries completely vandalize the benefits brought by high feature resolution and discrepancy measurement.
Phenomena above verify the rationality of analyzed mechanism in Section IV-C. Stereo matching in BEV takes effect and performs better when stringent requirements are met. Improper settings may reduce the RoadBEV-stereo to RoadBEV-mono.
@@
Image res. Voxel res. (cm) Volume constr. Abs err. (cm) RMSE (cm)
@@
$\checkmark$ $\checkmark$ $\checkmark$ 1.863 2.127
$\checkmark$ $\checkmark$ $\checkmark$ 1.960 2.263
$\checkmark$ $\checkmark$ $\checkmark$ 1.801 2.085
$\checkmark$ $\checkmark$ $\checkmark$ 0.730 0.970
$\checkmark$ $\checkmark$ $\checkmark$ 0.700 0.902
$\checkmark$ $\checkmark$ $\checkmark$ 1.851 2.122
$\checkmark$ $\checkmark$ $\checkmark$ 0.584 0.740
$\checkmark$ $\checkmark$ $\checkmark$ 0.563 0.697
$\checkmark$ $\checkmark$ $\checkmark$ 1.837 2.095

### v2

@@ Abstract
Road surface conditions, especially geometry profiles, enormously affect driving performance of autonomous vehicles. Vision-based online road reconstruction promisingly captures road information in advance. Existing solutions like monocular depth estimation and stereo matching suffer from modest performance. The recent technique of Bird’s-Eye-View (BEV) perception provides immense potential to more reliable and accurate reconstruction. This paper uniformly proposes two simple yet effective models for road elevation reconstruction in BEV named RoadBEV-mono and RoadBEV-stereo, which estimate road elevation with monocular and stereo images, respectively. The former directly fits elevation values based on voxel features queried from image view, while the latter efficiently recognizes road elevation patterns based on BEV volume representing discrepancy between left and right voxel features. Insightful analyses reveal their consistence and difference with perspective view. Experiments on real-world dataset verify the models’ effectiveness and superiority. Elevation errors of RoadBEV-mono and RoadBEV-stereo achieve 1.83 cm and 0.50 cm, respectively. The estimation performance improves by 50% in BEV based on monocular image. Our models are promising for practical applications, providing valuable references for vision-based BEV perception in autonomous driving. The code is released at https://github.com/ztsrxh/RoadBEV
@@ I Introduction
In recent years, the rapid development of unmanned ground vehicles (UGVs) has posed higher requirements for on-board perception systems. Real-time understanding of driving environment and conditions is vital for accurate motion planning and control [1, 2, 3]. For vehicles, road is the only contacting media with the physical world. Road surface conditions determine many vehicle characteristics and driving performance [4]. Road unevenness like bumps and potholes (as shown in Fig. 1 (a)) exacerbates ride experience of passenger vehicles, which is intuitively perceptible. Real-time road surface condition perception, especially the geometry elevation, immensely contributes to planning and control systems in enhancing safety and ride comfort [5, 6, 7].
@@ IV-C RoadBEV-stereo
Fig. 7 shows the architecture of RoadBEV-stereo. Similar with RoadBEV-mono and general stereo matching models, the left and right images first undergo feature extraction with shared weights. Settings about the CNN backbone are the same as RoadBEV-mono, i.e., EfficientNet-B6. The only difference is that the feature map resolution is interpolated to $\frac{1}{2}$ in the FPN. Explanations and analysis about resolution will be given below. The feature voxels are projected onto both the two image planes and query features from left and right perspectives. The left and right voxel features, $\boldsymbol{F}^{\prime}_{vox,l}$ and $\boldsymbol{F}^{\prime}_{vox,r}$, with the same shape $[C,N_{z},N_{y},N_{x}]$ are thereby obtained. In the architecture of stereo matching models, cost volume encoding similarities of the two feature maps is then built by correlation operation. We follow this paradigm and the voxel cost volume $\boldsymbol{V}_{discr}$ in BEV is derived as:
${\color[rgb]{0,0,0}\boldsymbol{V}_{discr}=\boldsymbol{F}^{\prime}_{vox,l}*\boldsymbol{F}^{\prime}_{vox,r}}$ (2)
where * is the point-wise multiplication. The volume construction here differs from stereo matching in terms of feature inputs. The vertical dimension of the feature voxel represents elevation candidates or proposals, which is equivalent to the disparity dimension in stereo matching. For stereo matching, input features are 3D and the disparity dimension is introduced during volume construction. Instead, our proposed method directly takes 4D voxel features as inputs whose proposal dimensions are introduced before volume construction.
Similar with RoadBEV-mono, the 4D BEV volume can be reshaped as a 3D tensor and processed with 2D convolutions, which efficiently saves computation. Nevertheless, to demonstrate its consistence with stereo matching models, we still treat it as a 4D volume and aggregate by 3D convolutions. Specifically, we stack six 3D-conv layers and three hourglass modules with $\frac{1}{2}$ resolution reduction. The channel dimension of BEV volume is finally reduced to one. Since the number of vertical voxels $N_{z}$ may be different from the number of elevation classes $N_{c}$, the elevation proposal dimension is interpolated to have $N_{c}$ channels. The elevation feature map $\boldsymbol{F}^{\prime}_{ele}\in\mathbb{R}^{N_{c}\times N_{y}\times N_{x}}$ is thereby derived. The following output layers are the same with RoadBEV-mono.
Figure 7: Architecture of RoadBEV-stereo. The voxel defined under the left camera’s coordinate queries pixel features of the left and right feature maps. We construct cost volume in BEV by correlation between left and right voxel features. 3D convolutions then aggregate the 4D volume in BEV.
@@
Despite the benefits, requirements for accurate RSR with stereo images in BEV are more strict. High-resolution feature map is necessary to cooperate with the precise but narrow search range. Road image features are similar in a certain area as road texture are repetitive without regular patterns. High-resolution feature is therefore vital for identifying slight feature variation. This is also the reason why we leverage image feature maps with $\frac{1}{2}$ resolution.
@@ V-A Implementation Details
We utilize the following metrics to evaluate model performance: absolute error (abs. err.), root mean squared error (RMSE), and ratio of grids with error greater than 0.5 cm (>0.5 cm). We implement the models on PyTorch platform and train on RTX 3090 GPUs. Batch size is 8 for all experiments. Learning rate is set as 8e-4 for RoadBEV-mono, while 5e-4 for RoadBEV-stereo. The optimizer is AdamW with weight decay 1e-4. The OneCycle learning rate scheduler with linear decreasing strategy is used. The images are cropped to 960*528 to meet the size requirements. The RoadBEV-mono is trained for 50 epochs, while 40 for RoadBEV-stereo. Pre-trained weights on ImageNet is loaded. The compared models in Table I are also trained for the same epochs, with their default configurations in the provided codes.
@@ V-B Performance and Comparison
Method Abs. err. (cm) RMSE (cm) >0.5 cm (%)
Mono LapDepth[45] 2.81 3.12 85.3
PixelFormer[46] 2.65 2.86 82.0
iDisc[47] 2.64 2.88 84.3
AdaBins[44] 2.59 2.79 82.4
RoadBEV(Ours) 1.83 2.07 78.6
Stereo IGEV-Stereo[48] 0.651 0.797 49.5
PSMNet[49] 0.654 0.785 50.1
CFNet[50] 0.647 0.760 50.8
ACVNet[51] 0.596 0.723 46.2
GwcNet[52] 0.588 0.711 44.9
DVANet[13] 0.546 0.685 40.9
RoadBEV (Ours) 0.503 0.609 37.0
Further, all the stereo-based models in Table I outperform the monocular-based with immense margins. Our RoadBEV-stereo also performs better than the others with noteworthy improvements. Although the absolute elevation error achieves 0.5 cm, only 37.0% grids have error bigger than 0.5 cm, indicating that most estimations are concentrated within a small error range. The error level of 5.0 mm covers almost all road unevenness causing vertical vibration to vehicles. Our model has great potential in on-board RSR for aiding planning and control systems of UGVs. The proposed BEV cost volume and estimation head jointly function and efficiently recognize feature patterns in BEV. The constraints on search range simplify similarity measurement and identification with effectiveness. Comparing the architectures of RoadBEV-mono (Fig. 5) and RoadBEV-stereo (Fig. 7), the principal promotion of RoadBEV-stereo lays in the utilization of extra information from another perspective (i.e., stereo image). The introduction of another view contributes to more than twice the performance enhancement. Therefore, stereo-based RSR is undoubtedly more promising and reliable than the monocular-based.
For more insights, we further visualize the distance-wise absolute error, as shown in Fig. 10. The whole grid is split into 15 segments along the longitudinal direction, with each segment representing a region of 33 cm long. Absolute elevation errors of grids inside the segments are averaged. Instead of a global but coarse metric, this analysis method digs into model performance and gives more comprehensive evaluation. Our RoadBEV-mono thoroughly outperforms the compared depth estimation models among the whole range with a significant margin. The superiority is more prominent at far distance comparing with LapDepth. As analyzed in Section IV-B. The AdaBins performs well as it defines the searching bins rather than direct fitting without effective constraints.
@@ V-D Ablation Studies for RoadBEV-mono
We first explore the influence of class interval on model performance, as listed in Table II. We set different class intervals between 0.4$\sim$2.0 cm, among which the metrics reach highest at 0.5 cm. Results also show that metric difference for class resolutions $\{0.5,0.8,1.0,1.6\}$ is slight. Potential reason is that we utilize soft argmin to regress elevation instead of picking the peak class. The weighted sum operation compensate accuracy decrease caused by discrete class interval. Over-high resolution at 0.4 cm results in $Nc$=100, which is too complicated to distinguish minute distinctness. Over-low resolution at 2.0 cm is too sparse to accurately represent the continuous elevation. Class interval is thereby set as 0.5 cm for both the proposed models.
@@
Class interval (cm) # classes Abs. err. (cm) RMSE (cm) >0.5 cm (%)
2.0 20 2.09 2.27 81.0
1.6 25 1.94 2.18 78.6
1.0 40 1.91 2.16 79.1
0.8 50 1.84 2.07 79.3
0.5 80 1.83 2.07 78.6
0.4 100 1.95 2.21 80.0
@@ V-E Ablation Studies for RoadBEV-stereo
The influence of volume construction, feature map resolution, and voxel interval are investigated. The first three experiments in Table IV exploit point-wise subtraction to construct BEV volume instead of correlation. The corresponding metrics are slightly lower than the correlation volume.
Results in the last six experiments imply that for a certain feature map resolution, distinct voxel intervals have very slight influence on estimation performance. The estimation head is effective for both sparse and dense feature samplings along the search ranges. The effect of feature map resolution is more significant, where the models with $\frac{1}{2}$ feature resolution are superior. For instance, the absolute error reduces from 0.550 cm to 0.518 cm when voxel resolution is 0.5 cm. Phenomena above verify the rationality of analyzed mechanism in Section IV-C. Stereo matching in BEV takes effect and performs better when high-resolution feature maps are given.
@@
Feature res. Voxel res. (cm) Volume constr. Abs err. (cm) RMSE (cm) >0.5 cm (%)
@@
$\checkmark$ $\checkmark$ $\checkmark$ 0.538 0.660 40.8
$\checkmark$ $\checkmark$ $\checkmark$ 0.512 0.623 38.2
$\checkmark$ $\checkmark$ $\checkmark$ 0.512 0.621 38.5
$\checkmark$ $\checkmark$ $\checkmark$ 0.550 0.683 41.3
$\checkmark$ $\checkmark$ $\checkmark$ 0.536 0.660 40.2
$\checkmark$ $\checkmark$ $\checkmark$ 0.535 0.655 40.4
$\checkmark$ $\checkmark$ $\checkmark$ 0.518 0.620 38.1
$\checkmark$ $\checkmark$ $\checkmark$ 0.503 0.609 37.0
$\checkmark$ $\checkmark$ $\checkmark$ 0.512 0.618 37.8